Shift report 20010819 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Aug 19 2001 - 00:32:11 EDT

  • Next message: Apache: "Shift report 20010819 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Aug 19 00:32:10 2001, Shift: 16:00-24:00
    Supervisor: YK Lee, with: A. Markeev
    Log book#7 pages 124 - 128
    
    Shift Summary:
    ------------------------------------------------------------
    1. 2 and 1/2 hours of good beem.
    2. Run 4278 has 72857 of Trigger 1 with polarity B.
       Two 15 minutes runs to test tiles as centrality trigger:
    	Run 4279 with Tile discrimination level at 0.75V
    	Run 4280 with Tile discrimination level at 0.50V
       Run 4281 is for drift velocity monitor.
       Run 4282 is a short run with magnet off.
            --Polarity change before run 4283---
    
       Run 4283 was started with Polarity A and is still
       running.
    3. T5 tripped once.
       T3 Anode 2 cannot be turned on. It has a large leakage
       (75 microamp) with anode off.
       About a dozen channels missing between channels 265 and
       295 for RICH ADC's.
       T1 Row 4 MaxADC per pad show dropped channels between
       31-50.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Aug 19 2001 - 00:33:01 EDT